There was a CD store. One day, they sold 80 CDs. The CDs either cost $15 or $12. If they sold $1104 worth of CDs, how many of each kind did they sell?
ALSO:
A store has $520 in one-dollar and five-dollar bills. If there are five times as many one-dollar bills as five-dollar bills, how many of each kind are there?

Found 2 solutions by rwm, marilynh:
Answer by rwm(914)

a+b=80
15a+12b=1104
a=48
b=32
1o+5f=520
o=5f
o=260
f=52

Answer by marilynh(7)

Let x = number of $15 CDs sold
Let y = number of $12 CDs sold

Since the total number of CDs sold is 80, x + y = 80.
Solve for one variable (x or y).
If we solve for x we get: x = 80 - y

The total amount sold is $1104, so 15x + 12y = 1104
Substitute x = 80 - y into the equation 15x + 12y = 1104
15(80 - y) + 12y = 1104
1200 - 15y + 12y = 1104
96 - 15y + 12y = 0
96 - 3y = 0
96 = 3y
32 = y
Which means there were 32 $12 CDs sold.

Substitute y = 32 in the equation x + y = 80
x + 32 = 80
x = 48
Which means there were 48 $15 CDs sold.
_____________________________________________________________________
Let x = the number of five dollar bills
Since there are five times as many $1 bills as $5 bills, 5x = number of one dollar bills.

$1 (5x) + $5 (x) = $520
5x + 5x = 520
10x = 520
x = 52
Which means there are 52 $5 bills
We can find out how many $1 bills there are in two ways:
1) Find out the value of the $5 bills: 52 x $5 = $260
The total of all the bills is $520, so $520 - $260 = $260
The value of the $1 bills is $260, so there are 260 $1 bills
2) We know there are five times as many $1 bills as $5 bills.
There are 52 $5 bills so there are 52 x 5 of the $1 bills
52 x 5 = 260